OpenStudy (anonymous):

The number of bacteria (b) in a batch of refrigerated food is given by the function b(t) = 20t2 − 70t + 300, where t is the temperature of the food in degrees Fahrenheit. The function t(h) = 2h + 3 gives the temperature of the food h hours after it is removed from refrigeration. Which expression correctly represents the number of bacteria (b) in the food as a function of the number of hours (h) the food is unrefrigerated?

OpenStudy (anonymous):

80h2 + 100h + 690 80h2 + 100h + 270 40h2 + 140h + 570 40h2 − 140h + 603

OpenStudy (trojanpoem):

Plug t(h) in b(t) so b(t(h) = 20( 2h + 3)^2 - 70 (2h+3) + 300 = 20 (4h^2 + 9 + 12h) - 140h - 210 + 300 80h^2 + 180 + 240h - 140 h - 210 + 300 = 80^2 + 270 + 100h = 80^2 + 100h + 270 If you found difficulty in understanding that, tag me.